OpenStudy (saadsword):

Please Help: If f(x - 1) = 2x + 3 for all values of x, what is the value of f(-3) ?

OpenStudy (saadsword):

@zepdrix

OpenStudy (saadsword):

@welshfella

OpenStudy (welshfella):

let f(x - 1) = a(x -1) + b = ax - a + b compare this with 2x + 3:- ax = 2x so a = 2 and -a + b = -2 + b = 3 so B = 3+2 = 5 so f(x) = 2x + 5

OpenStudy (welshfella):

now plug in x = -3 to find f)-3)

OpenStudy (welshfella):

* f(-3)

OpenStudy (saadsword):

i see so you have to find f(x) before f(-3)

OpenStudy (welshfella):

yeaea

OpenStudy (saadsword):

soo 2(-3) +5 = -1

OpenStudy (welshfella):

right

OpenStudy (saadsword):

thanks

OpenStudy (saadsword):

also there is another question that is bothering me

OpenStudy (welshfella):

ok but plz re post

OpenStudy (saadsword):

k